# Mixed Gauge-Global Symmetries, Elliptic Modes, and Black Hole Thermodynamics in Hořava-Lifshitz Gravity

### v1

@@ 6.1 Universal Horizons in general
where we can additionally note that $e(r)=(u\cdot\chi)^{2}-(s\cdot\chi)^{2}$, which ensures that the unit norm condition of the æther is satisfied Universal ; Jishnu . By recalling that $u$ is assumed to by hypersurface orthogonal, we should be able to easily extract the behavior of the surfaces of constant $T$ from (60) using $u=-NdT$. This requires us to note that $N\propto(u\cdot\chi)$, and is equal up to some function of the khronon. More precisely, $N=f(T)(u\cdot\chi)$. If we gauge fix $T$ via $\chi^{\alpha}\nabla_{\alpha}T=1$, we can see that Universal
@@ 9.1 Elliptic mode cosmic censorship as a motivator for universal horizons
While not the focus of this paper, there are two ways of interpreting the fact that slices must asymptotically approach the universal horizon. First, one can see that a given slice cannot cross the universal horizon in order to retain causality. Since the universal horizon is a leaf in the preferred foliation, another leaf crossing the universal horizon would generate a contradiction since the foliation, $\phi:\mathcal{M}\to\mathbb{R}\times S$, must be a bijection. The requirement that $\phi$ be a bijection prevents one point $p\in\mathcal{M}$ from being assigned two different æther times $T,T^{\prime}$ and ensures that $T$ is a well defined time coordinate that respects causality. A leaf in the foliation crossing the universal horizon (another leaf) would generate this exact contradiction of the injectivity of $T$.
9.2 Splittability
@@
Since Hořava Lifshitz gravity is a holographic theory of quantum gravity, the work in Harlow_Main ; Harlow_PRL immediately begs the question of how this mixed symmetry is allowed to exist. Since the mixed symmetry is not gauged along a slice $\Sigma$, and since there is a notion of entanglement wedges in Hořava Lifshitz gravity wedges , how does Hořava gravity avoid the proof given in Harlow_Main ; Harlow_PRL ? The answer is that the mixed symmetry is not a splittable symmetry: the assumption upon which the proof in AdS/CFT was built. This does not allow one to write the global symmetry operator $U(g,\partial\Sigma)$ of the boundary theory as a product of operators on subregions in the first place, so one can never construct a bulk field $\phi$ which will commute with bulk symmetry operators with support on entanglement wedges corresponding to strict subsets $R_{i}\subset\partial\Sigma$.

### v2

@@ 5.1 Universal Horizons in general
where we can additionally note that $e(r)=(u\cdot\chi)^{2}-(s\cdot\chi)^{2}$, which ensures that the unit norm condition of the æther is satisfied Universal ; Jishnu . By recalling that $u$ is assumed to be hypersurface orthogonal, we should be able to easily extract the behavior of the surfaces of constant $T$ from (60) using $u=-NdT$. This requires us to note that $N\propto(u\cdot\chi)$, and is equal up to some function of the khronon. More precisely, $N=f(T)(u\cdot\chi)$. If we gauge fix $T$ via $\chi^{\alpha}\nabla_{\alpha}T=1$, we can see that Universal
@@ 8 Remarks and future directions
8.1 No global symmetries in quantum gravity and splittability
8.1.1 Background
Now that we have seen how the mixed symmetry resolves outstanding issues in universal horizon thermodynamics we turn to a broader view on the existence of this symmetry in a putative consistent theory of quantum gravity in the first place. A key longstanding conjecture about what symmetries are allowed in quantum gravity is that there are no exact global symmetries in any complete and consistent theory of quantum gravity Banks_Constraints ; Banks ; Polchinski_Dualities ; Hsin:2020mfa ; Harlow_Main ; Harlow_PRL . This conjecture originally arose from deep considerations about black hole thermodynamics Hawking_particle and, as befitting the importance of symmetries, is intimately connected to other major conjectures about the structure of any viable quantum gravity theory including the weak gravity conjecture (c.f Harlow_WGC ), various completeness conjectures (c.f. Rudelius:2020orz ), and the distance conjecture (c.f. Cordova:2022rer ). Proofs that there are no global symmetries are also possible in more limited holographic constructions Harlow_PRL ; Harlow_Main , giving a clue that the conjecture is also relevant for holographic quantum gravity in general. As well, calculations in Euclidean quantum gravity indicate that global symmetries are not respected quantum mechanically Banks ; Strominger ; KLee_Wormhole ; ABBOTT1989687 ; COLEMAN1990387 .
Historically, the justification for the no global symmetries conjecture has been in the form of a heuristic argument regarding black holes that radiate via Hawking radiation Hawking_particle . More recently, the conjecture has been proven rigorously in AdS/CFT by Harlow and Ooguri in Harlow_Main ; Harlow_PRL using entanglement wedge reconstruction and splittability of boundary CFTs. In this section, we will review the heuristic justification for this conjecture and introduce the tension between the conjecture and Hořava-Lifshitz theory.
As we have seen above, Hořava-Lifshitz gravity possesses black holes which radiate and have corresponding first laws. As well, it is obvious that the theory is meant to be a consistent theory of quantum gravity. However, given the existence of the reparameterization symmetry outlined in Section 4, it is clear that there are symmetries which are not local gauge symmetries but rather mixed. In other words, for Hořava-Lifshitz gravity, the absence of global symmetries does not automatically imply that the symmetries are local gauge. Indeed as we have seen, the existence of a mixed symmetry is actually required to resolve outstanding problems with the thermodynamics. This may seem to generate the same contradiction that arose in the heuristic argument above. However, since the equation of motion governing the lapse $N$ is elliptic, as seen in (39), this seems to give a mechanism of generating the appropriate chemical potentials for the black holes to dissipate the elliptic charge and avoid the contradiction found in the heuristic argument. Furthermore, the promise for a full holographic description of Hořava-Lifshitz gravity with boundary Lifshitz field theories Horava_Lifshitz ; Horava_Membranes ; Iranian ; CFTno ; wedges begs the question of how it may avoid the existing holographic proofs that bulk symmetries must be long range gauge as found in Harlow_PRL ; Harlow_Main .
8.1.2 Splittability
@@
Since the mixed symmetry is not gauged along a slice $\Sigma$, and since there is a notion of entanglement wedges in Hořava Lifshitz gravity wedges , how does Hořava gravity avoid the proof given in Harlow_Main ; Harlow_PRL if Hořava gravity is holographic? The answer is that the mixed symmetry is not a splittable symmetry: the assumption upon which the proof in AdS/CFT was built. This does not allow one to write the global symmetry operator $U(g,\partial\Sigma)$ of the boundary theory as a product of operators on subregions in the first place, so one can never construct a bulk field $\phi$ which will commute with bulk symmetry operators with support on entanglement wedges corresponding to strict subsets $R_{i}\subset\partial\Sigma$.
There are two ways of interpreting the fact that slices must asymptotically approach the universal horizon. First, one can see that a given slice cannot cross the universal horizon in order to retain causality. Since the universal horizon is a leaf in the preferred foliation, another leaf crossing the universal horizon would generate a contradiction since the foliation, $\phi:\mathcal{M}\to\mathbb{R}\times S$, must be a bijection. The requirement that $\phi$ be a bijection prevents one point $p\in\mathcal{M}$ from being assigned two different æther times $T,T^{\prime}$ and ensures that $T$ is a well defined time coordinate that respects causality. A leaf in the foliation crossing the universal horizon (another leaf) would generate this exact contradiction of the injectivity of $\phi^{-1}$.